5floz. of a 2% alcohol solution was mixed with 11floz. of a 66% alcohol solution. Find the concentrationof the new mixture.
I need to know how to set the problem up! thanks

Answer by stanbon(75887)

5floz. of a 2% alcohol solution was mixed with 11floz. of a 66% alcohol solution. Find the concentration of the new mixture.
-----------------
5 fl oz of 2% has 0.02*5 = 0.10 oz active ingredient
11 fl oz of 66% has 7.26 oz of active ingredient
-----------------
A mixture of the two would give you 16 fl oz with 7.26+0.10 oz = 7.36 oz
of active ingredient or 7.36/16 = 46% concentration.
------------------------
Cheers,
Stan H.